OpenStudy (anonymous):

HEY HELP ME PLEASE LOL Explain the Product and Quotient Properties of Square Roots and include examples.

OpenStudy (hba):

NO NO NO NO NO

OpenStudy (anonymous):

product of two is simply like this.. x^m*x^n=x^(m+n) let is it is sqrtx into cuberoot x then =x^(1/2+1/3)=x^5/6

OpenStudy (anonymous):

Ayy be nice @hba

OpenStudy (anonymous):

And thanks :) @zaibali.qasmi

OpenStudy (anonymous):

I ned more info, please @zaibali.qasmi

OpenStudy (anonymous):

whhat you want more..?

OpenStudy (anonymous):

"Explain the Product and Quotient Properties of Square Roots and include examples." Did you do that or lol I dont know

OpenStudy (anonymous):

Product property of square roots: For all positive real numbers a and b, the square root of the product is the same as the product of the square roots: √a * √b = √ab example: √3 * √2 = √(3 * 2) = √6 Quotient property of square roots: For all positive real numbers a and b, b ≠ 0: the square root of the quotient is the same as the quotient of the square roots: √a / √b = √a/b example: √5 / √3 = √5/3

OpenStudy (anonymous):

get it..??

OpenStudy (hba):

Jk @cheyennestaggs

OpenStudy (anonymous):

lol @hba

OpenStudy (anonymous):

THANK YOU SO MUCH! :D @zaibali.qasmi

OpenStudy (anonymous):

its okay sweet...@cheyennestaggs